OpenStudy (anonymous):

Rewrite the equation in standard form and use your answer to fill in the three blanks below. (If one of the constants is negative, use a negative sign too.) The equation in standard form is

OpenStudy (texaschic101):

where is the original equation ?

OpenStudy (anonymous):

@texaschic101

OpenStudy (welshfella):

complete the square on x^2 + 6x:- take 1/2 of +6 and write it in the brackets:- x^2 + 6x = (x + 3^2 - 9 ( its -9 because (x+3)^2 = x^2 + 6x + 9)

OpenStudy (welshfella):

now do the same operation on y^2 - 4y

OpenStudy (welshfella):

then rearrange to get the numbers on right side and write the sum as a square